Find Vf at the given point. f(x,y,z) = x3 + y3 – 322 + z Inx, (1,5,5) Vf|(1,5,5) = Di+(\)j + ()k (Simplify your answers.)
Answer #1

= f (1,4, z) = (3x² + 2 ) ? + 3y2 h + (lnx-62) ř Ef (1,5,5) = ( 3 + 3) { + 3.52+ (ln(1) - 6.5) * Given frosy , 2) = x² + y² 3